# Pairwise Markov Chains for Volatility Forecasting

### v2

@@ 2.1 Main models for volatility forecasting
where $f$ is a model and $\epsilon_{t}$ is a random variable with mean equal to 00 and independent from the other random variables. We consider the random variable $Y_{t}$ taking its values in $\mathbb{R}^{2}$ such that the realization $y_{t}$ of $Y_{t}$ is $(\sigma^{2}_{t+1},u_{t}^{2})$. With each model presented in this subsection, the observed process $Y_{1:T}=(Y_{1},...,Y_{T})$ forms a Markov chain, and its probabilistic graph is represented in figure 1.
[…]
@@ 2.2.1 Definition
The Pairwise Markov Chain (PMC) [14] is a probabilistic graphical model that extends the well-known HMC [27, 28, 29, 30], improving its performance in many fields [17, 31, 32]. It considers the hidden random sequence $X_{1:T}=(X_{1},...,X_{T})$, where $X_{t}$ takes its value in $\Lambda_{X}$, and the observed sequence $Y_{1:T}=(Y_{1},...,Y_{T})$, where $Y_{t}$ takes its value in $\Omega_{Y}$. For this work, we consider $\Lambda_{X}=(\lambda_{1},...,\lambda_{N})$ as a finite discrete set, and $\Omega_{Y}=\mathbb{R}^{d}$, where $d\geq 1$. We also consider a homogeneous PMC, meaning that the laws defining it do not depend on $t$. The PMC defines the distribution of $(X_{1:T},Y_{1:T})$ as follows:
@@ 3.1 New algorithm
with $\text{with }\delta_{t+1}(x_{t+1})$ $=\sum\limits_{x_{t}}p(x_{t}|y_{1:t})\frac{p(x_{t+1}|x_{t})}{p(x_{t+1})}p(x_{t+1}|y_{t+1}).$
@@ 4 Results
In the context of our empirical study, we will use the GARCH(1, 1), the FNN(2), the FNN(3), and the FNN(2, 3) as base models. These models will be combined with PMCs having $2$, $3$, and $4$ hidden states. The model associating the PMC with $N$ hidden states and a model $f$ is denoted PMC($N$)-$f$.

### v3

@@ 2.1 Main models for volatility forecasting
where $f$ is a model and $\epsilon_{t}$ is a random variable with mean equal to 0 and independent from the other random variables. We consider the random variable $Y_{t}$ taking its values in $\mathbb{R}^{2}$ such that the realization $y_{t}$ of $Y_{t}$ is $(\sigma^{2}_{t+1},u_{t}^{2})$. With each model presented in this subsection, the observed process $Y_{1:T}=(Y_{1},...,Y_{T})$ forms a Markov chain, and its probabilistic graph is represented in figure 1.
$Y_{1}$$Y_{2}$$Y_{3}$$Y_{4}$
@@ 2.2.1 Definition
The Pairwise Markov Chain (PMC) [14] is a probabilistic graphical model that extends the well-known HMC [27, 28, 29, 30], improving its performance in many fields [17, 31, 32, 33]. It considers the hidden random sequence $X_{1:T}=(X_{1},...,X_{T})$, where $X_{t}$ takes its value in $\Lambda_{X}$, and the observed sequence $Y_{1:T}=(Y_{1},...,Y_{T})$, where $Y_{t}$ takes its value in $\Omega_{Y}$. For this work, we consider $\Lambda_{X}=(\lambda_{1},...,\lambda_{N})$ as a finite discrete set, and $\Omega_{Y}=\mathbb{R}^{d}$, where $d\geq 1$. We also consider a homogeneous PMC, meaning that the laws defining it do not depend on $t$. The PMC defines the distribution of $(X_{1:T},Y_{1:T})$ as follows:
@@ 3.1 New algorithm
$\text{with }\delta_{t+1}(x_{t+1})$ $=\sum\limits_{x_{t}}p(x_{t}|y_{1:t})\frac{p(x_{t+1}|x_{t})}{p(x_{t+1})}p(x_{t+1}|y_{t+1}).$
@@ 4 Results
In the context of our empirical study, we will use the GARCH(1, 1), the FNN(2), the FNN(3), and the FNN(2, 3) as base models. These models will be combined with PMCs having $2$, 33, and $4$ hidden states. The model associating the PMC with $N$ hidden states and a model $f$ is denoted PMC(NN)-$f$.